# BioX-Bridge: Model Bridging for Unsupervised Cross-Modal Knowledge Transfer across Biosignals

### v1

@@ Abstract
Biosignals offer valuable insights into the physiological states of the human body. Although biosignal modalities differ in functionality, signal fidelity, sensor comfort, and cost, they are often intercorrelated, reflecting the holistic and interconnected nature of human physiology. This opens up the possibility of performing the same tasks using alternative biosignal modalities, thereby improving the accessibility, usability, and adaptability of health monitoring systems. However, the limited availability of large labeled datasets presents challenges for training models tailored to specific tasks and modalities of interest. Unsupervised cross-modal knowledge transfer offers a promising solution by leveraging knowledge from an existing modality to support model training for a new modality. Existing methods are typically based on knowledge distillation, which requires running a teacher model alongside student model training, resulting in high computational and memory overhead. This challenge is further exacerbated by the recent development of foundation models that demonstrate superior performance and generalization across tasks at the cost of large model sizes. To this end, we explore a new framework for unsupervised cross-modal knowledge transfer of biosignals by training a lightweight bridge network to align the intermediate representations and enable information flow between foundation models and across modalities. Specifically, we introduce an efficient strategy for selecting alignment positions where the bridge should be constructed, along with a flexible prototype network as the bridge architecture. Extensive experiments across multiple biosignal modalities, tasks, and datasets show that BioX-Bridge reduces the number of trainable parameters by 88–99% while maintaining or even improving transfer performance compared to state-of-the-art methods.
@@ 3.3 Bridge Position Selection
$\operatorname*{arg\,min}_{m\in\{1,\dots,M\}}\frac{1}{|\mathcal{D}^{\text{(pair)}}|}\sum_{i=1}^{|\mathcal{D}^{\text{(pair)}}|}\mathcal{L}_{\text{probe}}\left(g_{\bm{\omega}}^{\text{(old)}}\left(\bm{h}_{m,i}^{\text{(new)}}\right),\hat{y}_{i}\right),$ (4)
where $\bm{h}_{m,i}=f^{\text{(new)}}_{\bm{\phi}_{\leq m}}(\mathbf{x}_{i}^{\text{(new)}})$ denotes the $i$-th sample’s intermediate representation from the $m$-th layer of the new modality model, and $\hat{y}_{i}=f^{\text{(old)}}_{\bm{\theta}}(\mathbf{x}_{i}^{\text{(old)}})$ denotes the pseudo label. $\mathcal{L}_{\text{probe}}$ denotes the empirical loss for the linear prober.
@@
$\operatorname*{arg\,min}_{l\in\{1,\dots,L\}}\operatorname{CKA}_{\text{linear}}\left(\bm{H}_{m}^{\text{(new)}},\bm{H}_{l}^{\text{(old)}}\right)$ (5)
@@ 3.4 Bridge Architecture
Models of different modalities operate in distinct representational spaces. The bridge network should be sufficiently parametrized to enable the projection and alignment of the two spaces. A naive bridge architecture is a full-rank linear layer, but this is prohibitively expensive because of the high-dimensional projection from the new modality to the old modality. For example, using LaBraM (Jiang et al., 2024) as $f^{(\text{new})}_{\bm{\phi}}$ and HuBERT-ECG (Coppola et al., 2024) as $f^{(\text{old})}_{\bm{\theta}}$, the projection would require $N^{\text{(new)}}_{m}\times d^{\text{(new)}}_{m}\times N^{\text{(old)}}_{l}\times d^{\text{(old)}}_{l}=181\times 200\times 93\times 512\approx 1.7$ billion parameters. To address the challenge of high-dimensional projection, we propose a prototype network. The prototype network consists of two modules, a prototype set, and a low-rank approximation module. Specifically, the prototype set, $\bm{P}\!\in\!\mathbb{R}^{N_{p}\times d^{\text{(old)}}_{l}}$, consisting of $N_{p}$ prototype vectors with embedding dimension $d^{\text{(new)}}_{m}$, introduces the flexibility to incorporate prior knowledge from $f^{(\text{old})}_{\bm{\theta}}$ by initializing the prototypes as tokens/feature maps from $\bm{{h}}^{\text{(old)}}_{l}$. The low-rank approximation module, consisting of $\bm{A}\!\in\!\mathbb{R}^{d^{\text{(new)}}_{m}\times r}$ and $\bm{B}\!\in\!\mathbb{R}^{r\times N^{\text{(old)}}_{l}N_{p}}$, reduces the number of trainable parameters through a low-rank factorization, while generating aggregation weights for prototype vectors, as illustrated in Figure 2.
@@ 3.5 Bridge Training
Naturally, the training objective for the bridge network is to align the intermediate representations in the $L$-th layer  [^2] Note: While alignment can be performed at any layer between the $l$-th and $L$-th layers, we empirically find that performing alignment at the $L$-th (i.e., final) layer yields better transfer performance.:
@@ 4.1 Experimental Setups
We consider the following datasets in the evaluation: (i) WESAD (Schmidt et al., 2018) is a wearable stress and affect detection dataset consisting of synchronized data from a wrist- and chest-worn device, collected from 15 subjects during a lab study. We select the ECG and PPG modality for a three-class (baseline/amusement/stress) classification task. (ii) FOG (Zhang et al., 2022) is a multimodal dataset for detecting freezing of gait in Parkinson’s Disease, collected from 12 patients. We select the EMG and EEG modality for a two-class (normal/FOG) classification task. (iii) ISRUC (Khalighi et al., 2016) is a sleep-staging dataset consisting of synchronized data from a polysomnography, collected from 118 subjects in a laboratory study. We select the EEG and ECG modality for a two-class (sleep/wake) classification task. Dataset splits and preprocessing are detailed in the appendix. Given the unbalanced nature of the datasets, we report Balanced Accuracy, F1-Weighted, and F1-Macro, following (Jiang et al., 2024; Pillai et al., 2025).
@@
For EEG, we adopt the base version of the LaBraM architecture with 5.8M parameters (Jiang et al., 2024). For ECG, we adopt the small version of the HuBERT-ECG architecture with 30.4M parameters (Coppola et al., 2024). For PPG, we adopt the small version of the PaPaGei architecture with 5.7M parameters (Pillai et al., 2025). For EMG, we adopt NormWear with 136.1M parameters (Luo et al., 2024). Note that LaBraM, HuBERT-ECG, and NormWear adopt a CNN-transformer architecture, while PaPaGei adopts a CNN architecture. All models are initialized with the pre-trained weights provided by the original publications. Note that biosignal foundation models are still early in their development, in comparison to foundation models for language and vision. While current models contain a relatively small number of parameters, our method for efficient cross-modal knowledge transfer would be even more valuable as they scale up.
@@
ISRUC
$\text{EEG (Old)}\rightarrow\text{ECG (New)}$
$\text{ECG (Old)}\rightarrow\text{EEG (New)}$
Methods
Input
BAcc $\uparrow$
F1-M $\uparrow$
F1-W $\uparrow$
Params $\downarrow$
Input
BAcc $\uparrow$
F1-M $\uparrow$
F1-W $\uparrow$
Params $\downarrow$
@@ Methods
FOG
$\text{EEG (Old)}\rightarrow\text{EMG (New)}$
$\text{EMG (Old)}\rightarrow\text{EEG (New)}$
Methods
Input
BAcc $\uparrow$
F1-M $\uparrow$
F1-W $\uparrow$
Params $\downarrow$
Input
BAcc $\uparrow$
F1-M $\uparrow$
F1-W $\uparrow$
Params $\downarrow$
@@
WESAD
$\text{ECG (Old)}\rightarrow\text{PPG (New)}$
$\text{PPG (Old)}\rightarrow\text{ECG (New)}$
Methods
Input
BAcc $\uparrow$
F1-M $\uparrow$
F1-W $\uparrow$
Params $\downarrow$
Input
BAcc $\uparrow$
F1-M $\uparrow$
F1-W $\uparrow$
Params $\downarrow$
@@ 4.2 Unsupervised Cross-Modal Knowledge Transfer Performance
We also observe that the knowledge transfer performance gap compared to Oracle varies across datasets and knowledge transfer directions. For example, on the ISRUC dataset, we observe approximately 20% balanced accuracy gap between BioX-Bridge (60.11%) and Oracle (80.13%) for EEG $\rightarrow$ ECG, but only 1% gap (62.55% vs. 63.54%) for ECG $\rightarrow$ EEG. From this, we can draw a few conclusions. (1) Based on Oracle’s results, the ISRUC task is more difficult for ECG (63.54%) than for EEG (80.13%), which is reasonable, as EEG has been considered the gold standard for sleep. (2) Our unsupervised training using BioX-Bridge can effectively produce an ECG model (60.11%) that is close to supervised ECG training (63.54%). We observe a 20% gap because the baselines and BioX-Bridge use the ECG as input, which is a less physiologically relevant modality for sleep than the EEG used by the Oracle. (3) Unsupervised cross-modal knowledge transfer performance using EEG (62.55%) is also constrained by the performance of the ECG teacher (63.54%), a consequence inherent to knowledge transfer itself. Therefore, our unsupervised training produces a relatively weaker EEG model (62.55%) compared to supervised EEG training (80.13%) since we are transferring from a weaker ECG model (63.54%).
@@ 4.3 Ablation Studies
We conduct ablation studies on the WESAD dataset and the direction of knowledge transfer $\text{(PPG}\rightarrow\text{ECG)}$. Additional results are presented in the appendix.
@@
To show that the bridge position selection strategy proposed in Section 3.3 is effective, we compare the unsupervised performance of cross-modal knowledge transfer at various positions in Table 2. For “Fixed”, we train nine bridges in predefined positions, which is a combination of the first, middle, and last layers of the old and new modality models. The results are the average over all nine positions in five seeds (see more results in the appendix).
@@ 5 Conclusion
To ensure reproducibility, we provide detailed descriptions of our experimental setups, including data preprocessing steps, model architectures, hyperparameters, and training procedures, in Section 4.1 and Appendix D. Our code implementation is available in the supplementary materials and will be available in a dedicated repository upon publication.
@@ Appendix B Limitations and Future Work
Although BioX-Bridge greatly reduces training computational requirements and improves the efficiency of cross-modal knowledge transfer, it depends on the availability of pre-trained models for each biosignal modality, an assumption that may not hold for emerging or underexplored biosignals. Furthermore, depending on the position of the bridge, the inference time of the bridged model could be longer. Finally, existing unsupervised cross-modal knowledge transfer frameworks rely on the availability of paired data, which is more difficult to collect than unimodal data. Future work focusing on transfer using unpaired data would extend applications to any combination of modalities.
@@ D.1.2 Dataset Split
Dataset split is summarized in Figure A1. The datasets contain synchronized data from the old and new modalities. We perform a subject-wise split for WESAD and ISRUC and sample-wise split for FOG (Zhang et al., 2022) to obtain four subsets $\mathcal{D}^{\text{(old)}},\mathcal{D}^{\text{(new)}},\mathcal{D}^{\text{(val)}}$, and $\mathcal{D}^{\text{(pair)}}$, at a ratio of 33%, 22%, 11%, and 33%, respectively. We use old modality data from $\mathcal{D}^{\text{(old)}}$ to train the linear prober $g_{\omega}^{\text{(old)}}$. New modality data from $\mathcal{D}^{\text{(new)}}$ is used to evaluate bridge performance in an unseen set. All data from $\mathcal{D}^{\text{(pair)}}$ and $\mathcal{D}^{\text{(val)}}$ are used to train and help select hyperparameters.
@@ D.1.3 BioX-Bridge and Baseline Implementation Details
For the Random baseline, we simulate a model that randomly assigns labels to samples uniformly across all classes. For CardioGAN, we adopt the pre-trained weights provided by the original publication. We do not perform any fine-tuning as CardioGAN has been trained with WESAD as one of its datasets. We adopt the preprocessing code provided by CargioGAN to prepare the PPG. For each recording, we generate the corresponding ECG using a sliding window of 4 seconds and a step size of 4 seconds. The generated ECG recording is then preprocessed following the HuBERT-ECG pipeline described in Section D.1.1.

### v2

@@ Abstract
Biosignals offer valuable insights into the physiological states of the human body. Although biosignal modalities differ in functionality, signal fidelity, sensor comfort, and cost, they are often intercorrelated, reflecting the holistic and interconnected nature of human physiology. This opens up the possibility of performing the same tasks using alternative biosignal modalities, thereby improving the accessibility, usability, and adaptability of health monitoring systems. However, the limited availability of large labeled datasets presents challenges for training models tailored to specific tasks and modalities of interest. Unsupervised cross-modal knowledge transfer offers a promising solution by leveraging knowledge from an existing modality to support model training for a new modality. Existing methods are typically based on knowledge distillation, which requires running a teacher model alongside student model training, resulting in high computational and memory overhead. This challenge is further exacerbated by the recent development of foundation models that demonstrate superior performance and generalization across tasks at the cost of large model sizes. To this end, we explore a new framework for unsupervised cross-modal knowledge transfer of biosignals by training a lightweight bridge network to align the intermediate representations and enable information flow between foundation models and across modalities. Specifically, we introduce an efficient strategy for selecting alignment positions where the bridge should be constructed, along with a flexible prototype network as the bridge architecture. Extensive experiments across multiple biosignal modalities, tasks, and datasets show that BioX-Bridge reduces the number of trainable parameters by 88–99% while maintaining or even improving transfer performance compared to state-of-the-art methods. Our code is available at: https://github.com/chenqi-li/BioX-Bridge.
@@ 3.3 Bridge Position Selection
$\operatorname*{arg\,min}_{m\in\{1,\dots,M\}}\frac{1}{|\mathcal{D}^{\text{(pair)}}|}\sum_{i=1}^{|\mathcal{D}^{\text{(pair)}}|}\mathcal{L}_{\text{probe}}\left(g_{\bm{\eta}}\left(\bm{h}_{m,i}^{\text{(new)}}\right),\hat{y}_{i}\right),$ (4)
where $\bm{h}_{m,i}^{\text{(new)}}=f^{\text{(new)}}_{\bm{\phi}_{\leq m}}(\mathbf{x}_{i}^{\text{(new)}})$ denotes the $i$-th sample’s intermediate representation from the $m$-th layer of the new modality model, and $\hat{y}_{i}=g_{\bm{\omega}}^{\text{(old)}}\circ f^{\text{(old)}}_{\bm{\theta}}(\mathbf{x}_{i}^{\text{(old)}})$ denotes the pseudo label. $\mathcal{L}_{\text{probe}}$ denotes the empirical loss for the linear prober $g_{\bm{\eta}}$. Note that the pseudo labels are simply the argmax of the teacher logits used in traditional knowledge distillation methods.
@@
$\operatorname*{arg\,max}_{l\in\{1,\dots,L\}}\operatorname{CKA}_{\text{linear}}\left(\bm{H}_{m}^{\text{(new)}},\bm{H}_{l}^{\text{(old)}}\right)$ (5)
@@ 3.4 Bridge Architecture
Models of different modalities operate in distinct representational spaces. The bridge network should be sufficiently parametrized to enable the projection and alignment of the two spaces. A naive bridge architecture is a full-rank linear layer, but this is prohibitively expensive because of the high-dimensional projection from the new modality to the old modality. For example, using LaBraM (Jiang et al., 2024) as $f^{(\text{new})}_{\bm{\phi}}$ and HuBERT-ECG (Coppola et al., 2024) as $f^{(\text{old})}_{\bm{\theta}}$, the projection would require $N^{\text{(new)}}_{m}\times d^{\text{(new)}}_{m}\times N^{\text{(old)}}_{l}\times d^{\text{(old)}}_{l}=181\times 200\times 93\times 512\approx 1.7$ billion parameters. To address the challenge of high-dimensional projection, we propose a prototype network. The prototype network consists of two modules, a prototype set, and a low-rank approximation module. The prototype set, $\bm{P}\!\in\!\mathbb{R}^{N_{p}\times d^{\text{(old)}}_{l}}$, consisting of $N_{p}$ learnable prototype vectors with embedding dimension $d^{\text{(old)}}_{l}$, introduces the flexibility to incorporate prior knowledge from $f^{(\text{old})}_{\bm{\theta}}$. To initialize the learnable prototype set, $\mathbf{P}\in\mathbb{R}^{N_{p}\times d^{(\mathrm{old})}_{l}}$, we randomly select $N_{p}$ token/feature maps from $\mathbf{h}^{(\mathrm{old})}_{l}$ of the training set samples. Specifically, $\mathbf{h}^{(\mathrm{old})}_{l}$ are the intermediate representations from the $l$-th layer of the old modality model. The low-rank approximation module, consisting of $\bm{A}\!\in\!\mathbb{R}^{d^{\text{(new)}}_{m}\times r}$ and $\bm{B}\!\in\!\mathbb{R}^{r\times N^{\text{(old)}}_{l}N_{p}}$, reduces the number of trainable parameters through a low-rank factorization, while generating aggregation weights for prototype vectors, as illustrated in Figure 2.
@@ 3.5 Bridge Training
Naturally, the training objective for the bridge network is to align the intermediate representations in the $L$-th layer  [^2] Note: While alignment can be performed at any layer between the $l$-th and $L$-th layers, we empirically find that performing alignment at the $L$-th (i.e., final) layer yields better transfer performance. We believe this is a result of error propagation. For alignment at the $l$-th layer, a small alignment error would grow as it propagates to the final layer. For alignment at the $L$-th layer, the error growth will be reflected in the alignment loss, enabling the bridge to take that into account and yield better downstream performance.:
@@ 4.1 Experimental Setups
We consider the following datasets in the evaluation: (i) WESAD (Schmidt et al., 2018) is a wearable stress and affect detection dataset consisting of synchronized data from a wrist- and chest-worn device, collected from 15 subjects during a lab study. We select the ECG and PPG modality for a three-class (baseline/amusement/stress) classification task. (ii) FOG (Zhang et al., 2022a) is a multimodal dataset for detecting freezing of gait in Parkinson’s Disease, collected from 12 patients. We select the EMG and EEG modality for a two-class (normal/FOG) classification task. (iii) ISRUC (Khalighi et al., 2016) is a sleep-staging dataset consisting of synchronized data from a polysomnography, collected from 118 subjects in a laboratory study. We select the EEG and ECG modality for a two-class (sleep/wake) classification task. Dataset splits and preprocessing are detailed in the appendix. Given the unbalanced nature of the datasets, we report Balanced Accuracy, F1-Weighted, and F1-Macro, following (Jiang et al., 2024; Pillai et al., 2025).
@@
For EEG, we adopt the base version of the LaBraM architecture with 5.8M parameters (Jiang et al., 2024). For ECG, we adopt the small version of the HuBERT-ECG architecture with 30.4M parameters (Coppola et al., 2024). For PPG, we adopt the small version of the PaPaGei architecture with 5.7M parameters (Pillai et al., 2025). For EMG, we adopt NormWear with 136.1M parameters (Luo et al., 2024). Note that LaBraM, HuBERT-ECG, and NormWear adopt a CNN-transformer architecture, while PaPaGei adopts a CNN architecture. All models are initialized with the pre-trained weights provided by the original publications. Note that biosignal foundation models are still early in their development, in comparison to foundation models for language and vision. Although current models contain a relatively small number of parameters, our method for efficient cross-modal knowledge transfer would be even more valuable as they scale up. As foundation models scale up, BioX-Bridge is efficient in terms of the bridge architecture to help reduce the number of trainable parameters. Furthermore, the bridge position selection strategy is lightweight in comparison to model training and scales linearly with the number of layers of the foundation models.
@@
ISRUC $\text{EEG (Old)}\rightarrow\text{ECG (New)}$ $\text{ECG (Old)}\rightarrow\text{EEG (New)}$
Methods Input BAcc $\uparrow$ F1-M $\uparrow$ F1-W $\uparrow$ Params $\downarrow$ Input BAcc $\uparrow$ F1-M $\uparrow$ F1-W $\uparrow$ Params $\downarrow$
@@
FOG $\text{EEG (Old)}\rightarrow\text{EMG (New)}$ $\text{EMG (Old)}\rightarrow\text{EEG (New)}$
Methods Input BAcc $\uparrow$ F1-M $\uparrow$ F1-W $\uparrow$ Params $\downarrow$ Input BAcc $\uparrow$ F1-M $\uparrow$ F1-W $\uparrow$ Params $\downarrow$
@@
WESAD $\text{ECG (Old)}\rightarrow\text{PPG (New)}$ $\text{PPG (Old)}\rightarrow\text{ECG (New)}$
Methods Input BAcc $\uparrow$ F1-M $\uparrow$ F1-W $\uparrow$ Params $\downarrow$ Input BAcc $\uparrow$ F1-M $\uparrow$ F1-W $\uparrow$ Params $\downarrow$
@@ 4.2 Unsupervised Cross-Modal Knowledge Transfer Performance
We also observe that the knowledge transfer performance gap compared to the oracles varies between datasets and knowledge transfer directions. For example, on the ISRUC dataset, we observe approximately 20% balanced accuracy gap between BioX-Bridge (60.11%) and the supervised EEG oracle (80.13%) for EEG $\rightarrow$ ECG, but only 1% gap between BioX-Bridge (62.55%) and the supervised ECG oracle (63.54%) for ECG $\rightarrow$ EEG. From this, we can draw a few observations: (1) Based on oracle results, the ISRUC task is more difficult for ECG (63.54%) than for EEG (80.13%), which is reasonable, as EEG has been considered the gold standard for sleep. (2) Our unsupervised training with BioX-Bridge can effectively produce an ECG model (60.11%) that performs similarly to the supervised ECG oracle (63.54%), despite the absence of labeled data. (3) There is a 20% gap between BioX-Bridge (60.11%) and the supervised EEG Oracle (80.13%) because the baselines and BioX-Bridge use the ECG as input, which is a less physiologically relevant modality for sleep than the EEG used by the supervised EEG oracle. (4) Unsupervised cross-modal knowledge transfer performance using EEG (62.55%) is also constrained by the performance of the ECG teacher (63.54%), a consequence inherent to knowledge transfer itself. Therefore, our unsupervised training produces a relatively weaker EEG model (62.55%) compared to supervised EEG training (80.13%) since we are transferring from a weaker ECG model (63.54%).
@@ 4.3 Ablation Studies
We conduct ablation studies on the WESAD dataset and the direction of knowledge transfer $\text{(PPG}\rightarrow\text{ECG)}$. Additional results are presented in Appendix D.6.
@@
To show that the bridge position selection strategy proposed in Section 3.3 is effective, we compare the unsupervised performance of cross-modal knowledge transfer at various positions in Table 2. For “Fixed”, we train nine bridges in predefined positions, which is a combination of the first, middle, and last layers of the old and new modality models. The results are the average over all nine positions in five seeds. A breakdown of the result at each of the nine predefined positions is available in Appendix D.5.
@@ 5 Conclusion
To ensure reproducibility, we provide detailed descriptions of our experimental setups, including data preprocessing steps, model architectures, hyperparameters, and training procedures, in Section 4.1 and Appendix D. Our code is available at: https://github.com/chenqi-li/BioX-Bridge.
Acknowledgements
The work of Chenqi Li was supported by the Cyril and Phillis Long Scholarship at The Queen’s College in partnership with the Clarendon Fund. The work of Tim Denison was funded by the Royal Academy of Engineering and supported by the NIHR Oxford Health Biomedical Research Centre (NIHR203316). The views expressed are those of the author(s) and not necessarily those of the NIHR or the Department of Health and Social Care. The Royal Academy of Engineering supported the work of Tingting Zhu under the Research Fellowship scheme.
@@ Appendix B Limitations and Future Work
Limitations: Although BioX-Bridge greatly reduces training computational requirements and improves the efficiency of cross-modal knowledge transfer, it depends on the availability of pre-trained models for each biosignal modality, an assumption that may not hold for emerging or underexplored biosignals. Furthermore, depending on the position of the bridge, the inference time of the bridged model could be longer.
Future Work: Whether the unsupervised cross-modal knowledge transfer method should be task-agnostic is an interesting next step to explore. A task-agnostic method would offer better generality, making it suitable for multi-task scenarios. However, such approaches can potentially limit task-specific performance compared to task-specific methods. Investigating other data-free and task-agnostic bridge position selection strategies would further improve the flexibility of BioX-Bridge for different application scenarios. In addition, existing unsupervised cross-modal knowledge transfer frameworks rely on the availability of paired data, which is more difficult to collect than unimodal data. Future work focusing on transfer using unpaired data would extend applications to any combination of modalities. Future works can also explore BioX-Bridge for datasets with more than two modalities and investigate how BioX-Bridge generalizes across datasets that share the same modalities, allowing a deeper understanding of modality interactions and dataset-specific adaptations (Zhang et al., 2022b; Liu et al., 2023; Fang et al., 2024).
@@ D.1.2 Dataset Split
Dataset split is summarized in Figure A1. The datasets contain synchronized data from the old and new modalities. We perform a subject-wise split for WESAD and ISRUC and sample-wise split for FOG (Zhang et al., 2022a) to obtain four subsets $\mathcal{D}^{\text{(old)}},\mathcal{D}^{\text{(new)}},\mathcal{D}^{\text{(val)}}$, and $\mathcal{D}^{\text{(pair)}}$, at a ratio of 33%, 22%, 11%, and 33%, respectively. We use old modality data from $\mathcal{D}^{\text{(old)}}$ to train the linear prober $g_{\omega}^{\text{(old)}}$. New modality data from $\mathcal{D}^{\text{(new)}}$ is used to evaluate bridge performance in an unseen set. All data from $\mathcal{D}^{\text{(pair)}}$ and $\mathcal{D}^{\text{(val)}}$ are used to train and help select hyperparameters.
@@ D.1.3 BioX-Bridge and Baseline Implementation Details
For the Random baseline, we simulate a model that randomly assigns labels to samples uniformly across all classes. For CardioGAN, we adopt the pre-trained weights provided by the original publication. We do not perform any fine-tuning as CardioGAN has been trained with WESAD as one of its datasets. We adopt the preprocessing code provided by CardioGAN to prepare the PPG. For each recording, we generate the corresponding ECG using a sliding window of 4 seconds and a step size of 4 seconds. The generated ECG recording is then preprocessed following the HuBERT-ECG pipeline described in Section D.1.1.